A car moving westward along a straight, level road increases its velocity uniformly from +19 m/s to +35 m/s in 10.2 s.How far did it move while accelerating? Answer in units of m.
1
Expert's answer
2011-02-02T08:11:44-0500
According to the equations of the motion
V=V0 + at
x = x0 +V0t +at2/2
Thus
a = (V-V0)/t = (35-19)/10.2 = 1.57 m/s2
x = 0 + 19*10.2 + 1.57*10.22/2 = 245.47 m.

Answer: 245.47 m.